OpenStudy (anonymous):

Let f(x) = 3x3 – 4x – 1 and g(x) = x + 1. Find f of x over g of x please I need to help asap!

OpenStudy (anonymous):

Let me see if I can help

OpenStudy (anonymous):

Okay, thank you

OpenStudy (anonymous):

This is a polynomial division problem, so we will take it step by step.

OpenStudy (anonymous):

Ahh, It's difficult because there is no easy way to type it up on here.

OpenStudy (anonymous):

hmm, let me see if I can find an easy way to type it up, and I'll try to get back to you.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

\[\text{Factor the polynomial using the rational roots theorem.} \] \[\frac{(x+1)(3x^2-3x-1)}{(x+1)}\\ \]\[\frac{\cancel{(x+1)}(3x^2-3x-1)}{\cancel{x+1} }\] \[=3x^2-3x-1 \]